OpenStudy (anonymous):

If x + y = -4, find the value of(I shall type in equation editor).

OpenStudy (anonymous):

\[x ^{3} + y ^{3} - 12xy + 64\]

OpenStudy (anonymous):

Concept involved: algebraic identities. I have racking my brains for the past half an hour. Someone please help.

OpenStudy (anonymous):

answer is zero

OpenStudy (anonymous):

Please explain...

OpenStudy (anonymous):

and the reason it is zero is because this factors as \[{(x+y+4) (x^2-x y-4 x+y^2-4 y+16)}\]

OpenStudy (anonymous):

Did you find it using identities?

OpenStudy (anonymous):

so if \(x+y=-4\) then the first term is 0

OpenStudy (anonymous):

OK, I get it but I don't understand how you factored to get whatever you got.

OpenStudy (anonymous):

Please explain the factoring part ...

OpenStudy (anonymous):

give me a minute

OpenStudy (anonymous):

ok ready?

OpenStudy (anonymous):

we want to make it look like \(a^3+b^3\) where \(a=x+y, b=4\) because this almost looks like \[(x+y)^3+64\] which you can factor as the sum of two cubes

OpenStudy (anonymous):

What do you do with -12xy?

OpenStudy (anonymous):

so \[(x+y)^3=x^3+y^3+3x^3y+3xy^2+y^3\] so if you replace \(x^3+y^3\) by \((x+y)^3\) you have to subtract off \(3x^2y+3xy^2\) and you get \[x ^{3} + y ^{3} - 12xy + 64=(x+y)^2+64-3x^2y-3xy^2-12xy\]

OpenStudy (anonymous):

let me know if that last step was clear. by replacing \[x^3+y^3\] by \[(x+y)^3\] i added \[3x^2y+3xy^2\] so i had to subtract it off

OpenStudy (anonymous):

Yeah I got that.

OpenStudy (anonymous):

i also made a typo it should have been \[x ^{3} + y ^{3} - 12xy + 64=(x+y)^3+64-3x^2y-3xy^2-12xy\]

OpenStudy (anonymous):

now you can factor the second part, the part after the 64 as \[-3x^2y-3xy^2-12xy=-3xy(x+y+4)\] easily

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now we can factor the whole thing, first the sum of two cubes \[a^3+b^3=(a+b)(a^2-ab+b^2)\] with \[a=x+y, b=4\] and it will work out

OpenStudy (anonymous):

you get \[(x+y)^3+64-3xy(x+y+4)\] \[=(x+y+4)((x+y)^2-4(x+y)+16)-3xy(x+y+4)\] giving you a common factor of \[x+y+4\] in both tersm

OpenStudy (anonymous):

yes and then factor the first part as the sum of two cubes as i wrote above. you see both terms have a common factor of \(x+y+4\)

OpenStudy (anonymous):

so you can rewrite as \[(x+y+4)((x+y)^2-4(x+y)+16-3xy)\] and you can do some algebra to clean up the second factor, but it is unnecessary for this problem since you already have the factor \(x+y+4\) that you want

OpenStudy (anonymous):

adding and subtracting to factor is a good trick to know, used also for completing the square. sometimes called "sophie germain" trick

OpenStudy (anonymous):

it is a bunch of algebra, but i think the steps i wrote are right. if it is unclear let me know

OpenStudy (anonymous):

Thank you so much!! ... I got it ... got to go now!

OpenStudy (anonymous):

it is also unclear how you were supposed to know this yw